# Optimizing Power Grid Topologies with Reinforcement Learning: A Survey of Methods and Challenges

### v1

@@ Abstract
Keywords Power grid reliability  $\cdot$ Topology Optimization  $\cdot$ Reinforcement Learning
@@ 1.1 Background
Electrical power grids form the backbone of modern society, responsible for transporting electricity from producers to consumers 24 hours a day, 365 days a year. Operating these grids is a demanding control task that requires continuous monitoring and frequent interventions by skilled experts to maintain network stability, keep power flow within the thermal limits of the equipment, and ensure voltage and frequency levels are met [1]. The increasing integration of renewable energy sources into the power grid and the rising energy demand add new layers of complexity to this task. With its unpredictable generation patterns, renewable energy makes power flows more variable and less predictable, altering how the network behaves and responds to disturbances. Due to the operational costs of simulation with the current models, operators often rely on their experience to resolve network problems in real-time. Continuing this manual approach may compromise system security or lead to significantly higher costs.
According to projections by [2], energy consumption in the Netherlands is expected to grow 30% between 2024 and 2033, with similar trends across Europe. Meanwhile, the share of controllable energy sources is expected to decrease by approximately 40% due to the shift towards renewable energy. As a result, Transmission System Operators (TSOs) will face greater challenges in maintaining network security and reliability, underscoring the need for new grid management strategies.
@@
Reinforcement learning (RL) has emerged as a promising approach that could assist network operators in decision-making, offering the potential to find cost-effective flexibilities that may currently go unnoticed by human operators. This is why, in 2019, the French TSO RTE launched the Learning to Run a Power Network (L2RPN) challenge [1], encouraging researchers to use RL for power network control (PNC). This challenge seeks to explore how RL can enhance real-time operations and assist with operational planning, ultimately supporting TSOs in navigating the growing complexities of modern power grids.
Reinforcement Learning Framework for Power Grid Management
To facilitate the L2RPN challenge and further research on this topic, RTE provides an open-source Python package Grid2Op [3] that simulates realistic power grids. The Grid2Op package enables researchers to develop a sequential decision-making RL model, termed agent, capable of controlling the power grid in real-time. The package provides environments with different episodes, each simulating a continuous sequence of power grid states at 5-minute intervals.
Grid2Op provides a typical setup of a Markov Decision Process (MDP) defined by a tuple $(\mathcal{S},\mathcal{A},p,r)$, where at each time step $t$ an agent observes a state $s_{t}\in\mathcal{S}$ from the environment and takes an action $a_{t}\in\mathcal{A}$. The Grid2Op environment computes and returns the next state $s_{t+1}\in\mathcal{S}$ to the agent with probability $p(s_{t+1}|s_{t},a_{t})$, which is the unknown state transition probability, after which the agent receives an immediate reward $r(s_{t},a_{t})\in\mathbb{R}$ according to a defined reward function. With this MDP formulation, reinforcement learning (RL) can be used to learn a (stochastic) policy $\pi(a_{t}|s_{t})$ that optimizes the expected discounted reward $\mathbb{E}_{\pi}[\sum_{t=0}^{T}\gamma^{t}r(s_{t},a_{t})]$, where $\gamma\in(0,1)$ is the discount factor.
In Grid2Op, the states $s_{t}\in S$ include all grid information such as the topology of the grid, which includes the connections of the electrical elements, the power flow on the lines, power generation, and load consumption. The topology of the power grid can be represented as a graph $G=(V,E)$, with nodes $v\in V$, as substations, and edges $e\in E$, as transmission lines [4]. Each substation connects various elements such as line ends, generators (producers), loads (consumers) and storage units (batteries, introduced in [5]).
@@
Within each substation, there are multiple busbars available to connect the elements. A consistent busbar connection makes a substation a singular node; differing connections split it, changing the local network topology. Grid2Op can compute the resulting power flow configurations from any topology change; see Fig. 1 for an example.
@@
Modifications in a substation’s busbar configuration are termed bus-splitting actions. By default, Grid2Op assigns two busbars per substation and only recently [^1] Starting from Grid2Op 1.9.9., it became possible to change this configuration and increase the number of busbars. Therefore, all competitions up to now and the solutions discussed in this paper consider only two busbars per substation. The total action space $\mathcal{A}$ consists of four types of actions:
@@
    Bus splitting;
@@
Rules of the game
To imitate the physical properties of the power network, each substation and each power line has a cooldown time, which is the time needed after activation before this element can be used adjusted again. Furthermore, there is a limited number of actions that can be done by an operator or by the technology, therefore, by default, Grid2Op can handle only one substation and one line-switching action in each time step.
Each power line has a thermal limit, which represents the maximum amount of power it can safely carry before it breaks. When a line is overloaded for too long, it will automatically be disconnected. When this happens, some time steps are needed before this element can be activated again.
In the context of the RL challenge, a game-over state will occur when an agent fails to safeguard the power grid, i.e., any network disconnect or isolation of a load or generator. The overarching objective in grid management is to control power flows in a way that prevents overloads and avoids cascading failures that could destabilize the entire network.
@@ 1.2 Research motivation and contributions
Since the introduction of the L2RPN challenge in 2019 [1], the field of RL applied to power grid topology optimization has seen remarkable growth. Several subsequent challenges have been organized [6, 7, 8, 5, 9], and the Grid2Op package has been extended to include larger, more complex networks. These updates introduce additional challenges such as unexpected line attacks, increased penetration of renewable energy, integration of battery storage, and sending timely alerts in case of dangerous contingencies. Each challenge has prompted innovative solutions from the research community, contributing to a rapidly evolving landscape.
@@
To address these challenges, this paper aims to provide a comprehensive and structured overview of the methods and challenges in RL applied to grid topology control. In doing so, our aim is to guide researchers in selecting effective strategies and identifying open research questions in this dynamic field.

### v2

@@ Abstract
Keywords Power grid reliability  $\cdot$ Topology Optimization  $\cdot$ Reinforcement Learning  $\cdot$ Learning to Run a Power Network
@@ 1.1 Background
Electrical power grids form the backbone of modern society, being responsible for transporting electricity from producers to consumers 24 hours a day, 365 days a year. Operating these grids is a demanding control task that requires continuous monitoring and frequent interventions by skilled experts to maintain network stability, keep power flow within the thermal limits of the equipment, and ensure voltage and frequency levels are met, as highlighted by [1]. The increasing integration of renewable energy sources into the power grid and the rising energy demand add new layers of complexity to this task. With its unpredictable generation patterns, renewable energy makes power flows more variable and less predictable, altering how the network behaves and responds to disturbances. Due to the operational costs of simulation with the current models, operators often rely on their experience to resolve network problems in real-time. Continuing this manual approach may compromise system security or lead to significantly higher costs. According to projections by [2], energy consumption in the Netherlands is expected to grow 30% between 2024 and 2033, with similar trends across Europe ([3]). Meanwhile, the share of controllable energy sources is expected to decrease by approximately 40% due to the shift towards renewable energy. As a result, Transmission System Operators (TSOs) will face greater challenges in maintaining network security and reliability, underscoring the need for new grid management strategies.
@@
With its recent advancements and breakthroughs ([4], [5], [6], [7] and [8]), Reinforcement learning (RL) has emerged as a promising approach that could assist network operators in decision-making, offering the potential to find cost-effective flexibilities that may currently go unnoticed by human operators. For this reason, in 2019, the French TSO RTE launched the Learning to Run a Power Network (L2RPN) challenge introduced in [1], encouraging researchers to use RL for power network control (PNC). This challenge seeks to explore how RL can enhance real-time operations and assist with operational planning, ultimately supporting TSOs in navigating the growing complexities of modern power grids. Moreover, the authors of [1] note that RL has already demonstrated potential in a variety of power system problems, reviewed in [9, 10], highlighting its suitability for complex control tasks in this domain.
1.2 Introduction to Power Grids
In this section, we provide a brief background on power grids to facilitate researchers from the RL community stepping into the field of PNC. For a more extensive introduction, see also [11].
1.2.1 From generation to consumption
Electrical power grids are typically divided into two layers, the transmission grid, controlled by TSOs, and the distribution grid, see Fig. 1. Power is supplied to the network by generators. In the past, these generators were often located close to the consumers, the loads, reducing the transmission distances of the power. However, generation sources such as hydro and nuclear power stations or wind and solar are often situated in geographically remote areas, far from load centers. To reduce resistive losses over these long distances, electricity in the transmission grid is transported at high voltage levels, between 100.000 - 765.000 volts. The voltage in homes in Europe and North America is typically 240 and 120 volts, respectively. Therefore, step-down transformers are used to reduce the voltage for the distribution grid. From the power network’s perspective, consumption is aggregated as load, and the grid operator’s responsibility generally ends at the step-down transformer. To allow for voltage conversion using transformers, the transmission grid operates under alternating current (AC). While AC allows for efficient voltage transformation, it also adds complexity compared to direct current (DC) systems, as power flows are governed by nonlinear equations involving both active and reactive power balances.
Figure 1: Schematic illustration of power distribution from generation sources to consumers.
A transmission power grid can be represented as a highly interconnected directed graph, with transmission lines represented as edges and substations as nodes, see Fig. 2 for a small example. Substations contain the transformers to step up and step down the voltages and connect the generation sources and loads to the grid.
Figure 2: Example of a transmission power grid represented as a directed graph. This example is part of the environment l2rpn_icaps_2021, available in Grid2Op (see Section 1.4).
Human operators continuously monitor and operate the power system from the TSO control centers, intervening when necessary. Operators must always respect three key constraints, described in [11]:
  • Keep the transmission lines within their thermal limits. Each transmission line has an upper limit to the amount of power that can flow through, determined by its materials’ thermal properties. Overloading can lead to overheating, causing lines to sag dangerously close to the ground or damaging critical equipment such as transformers and cables. In such cases, automatic or manual disconnection is often preferable to irreversible damage.
  • Maintain the voltage within a safe range.
  • Balance generation and load.
Electricity flows within a power network in the way prescribed by power flow physics. In particular, this means that power cannot be directed from a specific generator to a specific consumer. If a line is disconnected, the power originally flowing on that line redistributes itself across the remaining lines, potentially overloading others. The job of the network operator and the challenge of L2RPN is to efficiently manage the network while avoiding overloads and cascading failures.
Indirectly managing the power flows during congestion by controlling injections, such as redispatching power generation sources, is the most common and well-documented approach in the literature. However, this strategy imposes constraints on producers to be flexible, which comes with additional costs, as mentioned by [12]. TSOs can alleviate congestion also by adjusting the grid topology via two types of topological actions, which, in contrast, have a negligible cost. We distinguish two types of actions: line switching and bus splitting. Line switching consists of remotely disconnecting or reconnecting a transmission line. Another less explored option is bus splitting, modifications performed within a substation (using node breakers).
@@ 1.2.1 From generation to consumption
Within each substation, there are multiple busbars available to connect the elements. A consistent busbar connection makes a substation a singular node; differing busbar connections split the substation, changing the local network topology. Fig. 4 shows an example of how a bus-splitting action changes the power flows and relieves the congested line in a simulated environment of Grid2Op (the package provided for L2RPN discussed in Section 1.4).
1.2.2 Topology Optimization
The challenges and solutions of L2RPN discussed in this survey paper mostly focus on topology optimization strategies, i.e., those using discrete actions such as line switching and substation reconfigurations. Continuous actions, such as redispatching, are only considered in a few solutions, often as a secondary option.
Topology optimization is a promising tool for relieving congestion, as it exploits existing infrastructure without resorting to costly interventions such as redispatching, load shedding, or demand response, [13, 1]. Despite this potential, topology optimization remains underutilized. Traditionally, the power systems community relies on optimization approaches to solve optimal power flow (OPF) problems, including augmented OPF with discrete topological actions, [14, 15, 16]. However, existing optimization techniques struggle to scale up, due to the problem’s nonlinear and discrete combinatorial nature, leading to a large and complex optimization space. Including all topology configuration actions results in a nonconvex mixed-integer nonlinear programming problem, which is NP hard. Most of the topology optimization research focuses on line switching [17, 18, 19, 20, 21, 22], but neglects the more complex bus-splitting actions, which could offer much more flexibility to the operator [12]. While bus-splitting is explored in [23, 24, 25, 26], many of these methods still exhibit one or more of the following limitations:
  • they are computationally expensive,
  • they are limited to only a small number of substations with bus-splitted options,
  • they rely on linear (DC) approximation, which neglects reactive power and voltage constraints, and/or
  • they treat the problem as a static single snapshot, whereas power network control is inherently a sequential decision problem that evolves over time [27].
These shortcomings highlight the need for new approaches, where RL, a fast-developing field of research specifically designed for sequential decision problems, could play a pivotal role.
1.3 Introduction to Reinforcement Learning
This section provides a brief background on reinforcement learning (RL) for researchers in the power systems community interested in its application to power network control (PNC). For a comprehensive overview of RL, we refer to the book by [28].
1.3.1 Key concepts reinforcement learning
Reinforcement learning is a machine learning technique where an agent (the learner and decision maker) learns from experience how to make sequential decisions in an environment, guided by a reward (feedback).
The sequential decision problem is typically modeled as a Markov Decision Process (MDP), which consists of:
  • State space $\mathcal{S}$. A state $s_{t}\in\mathcal{S}$ is the snapshot of the environment observed by the agent at time $t$, usually represented by a vector or matrix.
  • Action space $\mathcal{A}$ (or $\mathcal{A}_{s}$ the set of available actions from state $s$). After observing the state, the agent interacts with the environment at time $t$ by taking an action (decision) $a_{t}\in\mathcal{A}$ according to its policy $\pi$.
  • Transition function $p:\mathcal{S}\times\mathcal{S}\times\mathcal{A}\to[0,1]$. The quantity $p(s_{t+1}|s_{t},a_{t})$ describes the probability that action $a_{t}$ in state $s_{t}$ leads to state $s_{t+1}$.
  • Reward function $r:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$. The quantity $r(s_{t},a_{t},s_{t+1})$ represents the reward received after transitioning from $s_{t}$ to $s_{t+1}$ via action $a_{t}$. In some settings, this reduces to $r(s_{t},a_{t})$ or $r(s_{t})$.
  • Discount factor $\gamma\in(0,1)$, which in a discounted MDP setting, determines the relative importance of future rewards.
Section 1.4 presents the specific MDP setup for the L2RPN problem.
The RL agent interacts with the environment according to a (stochastic) policy $\pi:\mathcal{S}\times\mathcal{A}\to[0,1]$, where $\pi(a_{t}|s_{t})$ defines the probability of choosing the action $a_{t}$ in state $s_{t}$. The goal in RL is for the agent to learn a policy $\pi$ that maximizes the expected discounted return:
$V_{\pi}(s):=$ $\mathbb{E}_{\pi}\left[\sum_{t=0}^{T}\gamma^{t}r(s_{t},a_{t})\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ s_{0}=s\right].$
The function $V_{\pi}$ is also referred to as the state-value function under policy $\pi$. A fundamental property of RL introduced by [29] is the recursive property of the value function, which can be stated as
$V_{\pi}(s)=$ $\sum_{a\in\mathcal{A}}\pi(a|s)\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s,a)\left[r(s,a)+\gamma V_{\pi}(s^{\prime})\right]$
$=$ $\mathbb{E}_{\begin{subarray}{c}a\sim\pi\\ s^{\prime}\sim P\end{subarray}}\left[r(s,a)+\gamma V_{\pi}(s^{\prime})\right].$
where $s^{\prime}\sim P$ is shorthand for $s^{\prime}\sim p(\cdot|s,a)$.
Similarly, the action-value function (or $Q$-value function) for policy $\pi$ is defined by
$Q_{\pi}(s,a):=$ $\mathbb{E}_{\pi}\left[\sum_{t=0}^{T}\gamma^{t}r(s_{t},a_{t})\middle|s_{0}=s,a_{0}=a\right]=\mathbb{E}_{s^{\prime}\sim P}\left[r(s,a)+\gamma V_{\pi}(s^{\prime})\right].$
A policy $\pi^{*}$ that maximizes the expected return is an optimal policy. The optimal state-value function defined by $V^{*}(s):=\max_{\pi}V_{\pi}(s)$, satisfies the Bellman optimality equation
$V^{*}(s):=\max_{a\in\mathcal{A}}\mathbb{E}_{s^{\prime}\sim P}\left[r(s,a)+\gamma V^{*}(s^{\prime})\right].$
In terms of the $Q$-values, it can be equivalently rephrased as
$Q^{*}(s,a):=$ $\mathbb{E}_{s^{\prime}\sim P}\left[r(s,a)+\gamma\max_{a^{\prime}\in\mathcal{A}}Q^{*}(s^{\prime},a^{\prime})\right].$
The general idea of RL is to collect data by interacting with the environment and use this to find an approximation of the value function $V$ or $Q$ and/or to update the policy $\pi$.
Deep RL
The algorithms applied to the L2RPN problem usually make use of the so-called Deep RL. In Deep RL (DRL), the value functions and policies are approximated using a neural network, which is referred to as a parameterized policy and value function, where the parameters, often denoted by $\theta$ or $\phi$, represent the weights and biases of the neural network. This function approximation enables scalability to large or continuous state/action spaces, in contrast to table-based methods like tabular $Q$-learning.
1.3.2 RL methods and terminology
This survey discusses a variety of (advanced) RL algorithms applied to L2RPN. Describing each algorithm is outside the scope of this survey; here, we provide a few references and highlight the key distinctions between the algorithms discussed in this paper. To help the reader, we categorize the main algorithms used in L2RPN in Footnote 1.
Figure 3: An overview of the algorithms applied to L2RPN categorized by the features discussed in this section. [^1] This figure is inspired by https://spinningup.openai.com/en/latest/spinningup/rl_intro2.html.
Model-based vs. model-free
One of the main distinctions made in RL is between methods that require a model of the environment and methods that do not.
Model-based methods use a model of the environment (either known or learned) to simulate future outcomes and plan ahead. However, an accurate model of the environment is often unavailable or the model needs to be learned, which comes with additional challenges. The most famous example of a model-based approach is AlphaZero by [7], also applied in L2RPN.
Model-free methods learn directly from experience without requiring a model. These methods tend to be easier to implement and tune, making them a more popular choice in L2RPN competitions and most other research. Learning is typically done using some variant of temporal-difference (TD) updates. A basic TD method update of the value function is
$V_{k+1}(s_{t})\leftarrow V_{k}(s_{t})+\alpha\left[R_{t+1}+\gamma V_{k}(s_{t+1})-V_{k}(s_{t})\right],$
where $R_{t+1}$ and $s_{t+1}$ are the observed reward and state at time $t+1$ and $\alpha\in[0,1]$ is the learning rate. The quantity $V_{k}$ is the old estimate of the value function, while $V_{k+1}$ is the new estimate in iteration $k$. The quantity $R_{t+1}+\gamma V_{k}(s_{t+1})$ is the TD target.
Value-based vs. policy-based methods
Another distinction made between RL methods is whether they are value-based, policy-based, or a hybrid version of the two, referred to as an actor-critic method.
In value-based methods, the aim is to learn the values of the states or state-action pairs and then derive a policy from these. A breakthrough in RL was the development of $Q$-learning, an off-policy value-based TD method. In the context of L2RPN, some approaches use a DRL version of $Q$-learning (DQN), i.e., Double-DQN (DDQN) and Dueling-Double-DQN (D3QN), by [4] and [5].
Policy-based methods learn a policy directly rather than searching for a value function and extracting a policy. This is particularly well-suited for dealing with large or continuous action spaces. In Deep RL, they typically rely on policy gradient methods, which optimize the parameterized policy $\pi_{\theta}(a|s)$ by taking the gradient ascent of $J$, the expected return of a policy $\pi_{\theta}$, with respect to $\theta$, where
$J(\pi_{\theta}):=V_{\pi_{\theta}}(s_{0})=\mathbb{E}_{\pi_{\theta}}\left[R_{\tau}\right]$
and $\tau=(s_{0},a_{0},s_{1},a_{1},\dots)$ is a possible trajectory or episode following policy $\pi_{\theta}$. Typically, policy gradient methods are on-policy and learn stochastic policies, meaning that the policy specifies a probability distribution over the actions. Proximal Policy Optimization (PPO) by [6] and Cross Entropy Method (CEM) by [30] are policy-based methods used in the context of L2RPN. However, CEM uses population-based search and does not rely on policy gradients.
Actor-critic methods combine both paradigms, using a learned value function to reduce the variance of policy gradient updates. Soft Actor-Critic (SAC) and (Asynchronous) Advantage Actor-critic (A2C/A3C) are examples of actor-critic methods applied in L2RPN.
On-policy vs. off-policy
On-policy methods learn from data generated by the current policy being improved. In contrast, off-policy methods can learn from experience generated by a different behavior policy, for example $\epsilon$-greedy with respect to current $Q$-values, enabling more efficient data reuse and exploration. Using a behavior policy different from the current policy allows for applying techniques such as prioritized experience replay and different exploration strategies, as discussed in Sections 3.1.7 and 3.1.8.
1.4 Problem Description
This section describes the platform provided for L2RPN and how the PNC problem is formulated as a Markov Decision Process in Section 1.4.1. Additionally, Section 1.4.2 describes the rules within the L2RPN environment.
1.4.1 Reinforcement Learning Framework for Power Grid Management
To facilitate the L2RPN challenge and further research on this topic, RTE provides an open-source Python package Grid2Op ([31]) that simulates realistic power grids. The Grid2Op package enables researchers to develop a sequential decision-making RL model, termed agent, capable of controlling the power grid in real-time. The package provides environments with different episodes, each simulating a continuous sequence of power grid states at 5-minute intervals.
Grid2Op provides a typical setup of a Markov Decision Process (MDP) defined by a tuple $(\mathcal{S},\mathcal{A},p,r)$, where at each time step $t$ an agent observes a state $s_{t}\in\mathcal{S}$ from the environment and takes an action $a_{t}\in\mathcal{A}$. The Grid2Op environment computes and returns the next state $s_{t+1}\in\mathcal{S}$ to the agent with probability $p(s_{t+1}|s_{t},a_{t})$, which is the unknown state transition probability, after which the agent receives an immediate reward $r(s_{t},a_{t})\in\mathbb{R}$ according to a defined reward function. This agent-environment interaction is visualized in Fig. 4. With this MDP formulation, reinforcement learning (RL) can be used to learn a (stochastic) policy $\pi(a_{t}|s_{t})$ that optimizes the expected discounted reward $\mathbb{E}_{\pi}[\sum_{t=0}^{T}\gamma^{t}r(s_{t},a_{t})]$, where $\gamma\in(0,1)$ is the discount factor.
@@ 1.4.1 Reinforcement Learning Framework for Power Grid Management
In Grid2Op, the states $s_{t}\in S$ include all grid information such as the topology of the grid, which includes the connections of the electrical elements, the power flow on the lines, power generation, and load consumption. As mentioned in Section 1.2, the topology of the power grid can be represented as a graph, which we formalize as graph $G=(V,E)$, with nodes $v\in V$, as substations, and edges $e\in E$, as transmission lines, as described in [32]. Each substation connects various elements such as line ends, generators (producers), loads (consumers) and storage units (batteries, introduced in [33]).
The total action space $\mathcal{A}$ consists of four types of actions:
@@
    Bus splitting [^2] By default, Grid2Op assigns two busbars per substation and only recently, starting from Grid2Op 1.9.9., it became possible to change this configuration and increase the number of busbars. Therefore, all competitions up to now and the solutions discussed in this paper consider only two busbars per substation., shown in Fig. 4 and explained in Section 1.2.1;
@@
1.4.2 Rules of the game
As discussed in Section 1.2.1 grid operators need to keep the transmission lines within their thermal limit, maintain the voltage within a safe range and balance the generation and load. Similarly, agents in the L2RPN context must ensure that no line exceeds its thermal limit. Line loading, denoted by $\rho$, is defined by the ratio between the amount of power flow passing through and the thermal limit. When a line $e\in E$ is overloaded for too long, i.e. $\rho_{e}>1.0$, it will automatically disconnect and a specific number of time steps is needed before this element can be activated again.
Voltage control, and generation-load balance are handled by the environment in the current L2RPN challenge. However, redispatching generation can help relieve line congestion and can be considered a useful action in the RL challenge.
To mimic the physical properties of the power network, each substation and each power line has a cooldown time, which is the time needed after activation before this element can be adjusted again. Furthermore, there is a limited number of actions that can be done by an operator or by the technology, therefore, by default, Grid2Op can handle only one substation and one line-switching action in each time step.
In the context of this RL challenge, a game-over state will occur when an agent fails to safeguard the power grid, i.e., any network disconnect or isolation of a load or generator. The overarching objective in grid management is to control power flows in a way that prevents overloads and avoids cascading failures that could destabilize the entire network.
The rules described above align with an $N-0$ setting, meaning no contingencies are anticipated. By adding an adversarial agent, or opponent, in the L2RPN challenge described in Section 2.3, the proposed RL solutions are required to act more securely. In real-world operations, TSOs typically adhere to the stricter $N-1$ reliability standards, ensuring that the grid remains secure even after the failure of any single grid component.
@@ 1.5 Research motivation and contributions
Since the introduction of the L2RPN challenge in 2019 in [1], the field of RL applied to power grid control has seen remarkable growth. Several subsequent challenges have been organized ([11, 34, 35, 33, 36]) and the Grid2Op package has been extended to include larger, more complex networks. These updates introduce additional challenges such as unexpected line attacks, increased penetration of renewable energy, integration of battery storage, and sending timely alerts in case of dangerous contingencies. Each challenge has prompted innovative solutions from the research community, contributing to a rapidly evolving landscape.
@@
To address these challenges, this survey paper aims to provide a comprehensive and structured overview of the methods and challenges in reinforcement learning (RL) applied to grid topology control. Our aim is to guide researchers in selecting effective strategies and more easily identifying open research questions in this fast-moving field.